Jim's age is 4 years less than 3 times Fred's age, if Fred's age is doubled and added to Jim's age the result is 56. How old are Jim and Fred?
Answer by ptaylor(2198)

Let x=Fred's age
Then Jim's age = 3x-4
If we double Fred's age, we get 2x
Now if we add this to Jim's age we get:
2x+(3x-4)and we are told that this equals 56, so our equation to solve is:
2x+3x-4=56 add 4 to each side
2x+3x-4+4=56+4 collect like terms
5x=60 divide each term by 5
x=12-----------------------------------------------Fred's age
3x-4=3*12-4=36-4=32 ------------------------------Jim's age
CK
Jim's age is 4 years less than 3 times Fred's age
3*12-4=32
32=32
if Fred's age is doubled and added to Jim's age the result is 56.
2*12+32=56
56=56

Hope this helps----ptaylor